# Interstellar medium gas heating by primordial black holes and dark matter particles

### v1

@@ I Introduction
Dark matter (DM), the enigmatic source of gravity in the universe, remains one of the fundamental challenges in particle physics and cosmology Arbey2021PPNP . Among the most compelling candidates for DM is the primordial black hole (PBH) Carr1975APJ ; Carr2020ARNPS ; Carr2021RPP ; Auffinger2023PPNP , which formed during the early universe. While PBHs lose mass through the emission of particles via Hawking radiation Hawking1975CMP , those with an initial mass exceeding $10^{15}$ g can survive in the current universe without fully evaporating. Although it is unlikely that all DM is composed solely of PBHs across the broad mass spectrum Carr2021RPP , the possibility of PBH formation in numerous early universe models suggest it is reasonable to consider at least a portion of DM consists of PBHs.
@@
If only a portion of DM consists of PBHs, what constitutes the remainder? Since DM has eluded detection through electromagnetic observations, it is typically assumed to lack electromagnetic interactions. Under this assumption, electrically neutral, massive, and stable particles predicted by theories beyond the standard model are considered promising DM candidates Arbey2021PPNP . However, some particle DM models propose weak electromagnetic interactions. For example, dark photons may couple very weakly to electrically charged particles via kinetic mixing with ordinary photon Dubovsky2015JCAP . If dark photons is massive, they could serve as DM. Millicharged particles are another potential DM candidate Holdom1986PLB . Additionally, the temperature at cosmic dawn being only half of the expected value Bowman2018Nature can be explained by a Coulomb-like interaction between millicharged particles and neutral hydrogen (${\rm H_{I}}$) Barkana2018Nature .
@@
In recent years, the Leo T dwarf galaxy IrwinAPJ2007 , along with other astrophysical gas clouds, has been used to study the heating of interstellar medium gas by PBHs and DM particles. Leo T, located in the Leo constellation, is classified as a transitional object between dwarf spheroidal and dwarf irregular galaxies. Unlike dwarf galaxies within the Milky Way’s virial radius, which often lack gas due to ram pressure stripping Gunn1972APJ ; Grcevich2010APJ , Leo T, being beyond this radius, can retain significant amounts of gas. When DMs (such as PBHs, dark photons, or millicharged particles) heat the interstellar medium, the heating rate $Q$ must be equal or less than the astrophysical cooling rate $C$, $Q\leq C$. If $Q$ exceeds $C$, the temperature of the interstellar medium gas would increase, potentially altering the ionization ratio of the gas Kim2021MNRAS ; Lu2021APJL ; Laha2021PLB ; Wadekar2021PRD ; Takhistov2022JCAP ; Wadekar2023PRD ; Melikhov2023PRD ; Dubovsky2015JCAP . In the literature, the heating rate and the cooling rate are depicted as $\dot{Q}$ and $\dot{C}$, respectively; however, we omit ‘dot’ in this paper.
@@
Previously, various sources have been considered for heating interstellar medium gas, including PBHs via Hawking radiation Kim2021MNRAS ; Laha2021PLB ; Melikhov2023PRD , PBHs via accretion disks Lu2021APJL ; Takhistov2022JCAP ; Wadekar2023PRD , compact DM Wadekar2023PRD , dark photons (hidden photons) Dubovsky2015JCAP ; Bhoonah2019PRD ; Bhoonah2021PRD ; Wadekar2021PRD ; Prabhu2023PRD , millicharged particles Bhoonah2018PRL ; Wadekar2021PRD , vector portal DM Bhoonah2019PRD , Higgs portal DM Wadekar2022PRD , sterile neutrinos Wadekar2022PRD , dark baryons Wadekar2022PRD , axions and axion like particles Wadekar2023PRD ; Wadekar2022PRD , excited DM state Wadekar2022PRD , magnetically charged DM Wadekar2023PRD , composite DM Bhoonah2021PRD ; Wadekar2023PRD , annihilating DM Chen2023MNRAS , electromagnetic dipole moment of DM Bi2023EPJC , and DM-baryon interactions Bhoonah2019PRD ; Shoji2024MNRAS . These studies assume that only one type of DM is responsible for heating the interstellar medium gas. For example, in Ref Kim2021MNRAS ; Laha2021PLB ; Melikhov2023PRD ; Lu2021APJL ; Takhistov2022JCAP , PBHs are considered to be part of the DM and responsible for heating the gas, while other DM are not. Similarly, Ref. Bhoonah2018PRL assumes that millicharged particles are the sole contributors to the heating of the gas, with other DM having no effect.
@@
In this study, we consider DM to consist of two components: PBHs and additional DM particles. Unlike previous studies Kim2021MNRAS ; Laha2021PLB ; Melikhov2023PRD ; Lu2021APJL ; Takhistov2022JCAP , which assume that only one type of DM particles heats the gas, we propose that both PBHs and DM particles contribute to the heating of the gas simultaneously in the Leo T dwarf galaxy. We specifically consider dark photons or millichaged particles as the DM particles. The heating mechanism for PBHs is either through Hawking radiation or heat transport from the accretion disk. In contrast, the DM particles heat the gas via mechanism specific to their particle models.
@@
The remainder of this article is organized as follows. Section II provides a brief review of the Leo T profile and the current constraints on PBHs. Before exploring the simultaneous heating of the gas in Leo T by both PBHs and DM particles, we first estimate the heating contribution from individual sources, including light PBHs, heavy PBHs, dark photons, and millichaged particles. The main discussion in section II are based on Ref. Laha2021PLB for light PBHs, Ref.Lu2021APJL ; Takhistov2022JCAP for heavy PBHs, and Ref. Wadekar2021PRD for dark photons and millicharged particles. In section III, we assess the combined effect of PBHs and DM particles (dark photons or millicharged particles) on gas heating. We discuss constraints on the PBH fraction, the kinetic mixing parameter of dark photons, and the charge parameter for millicharged particles. The article concludes with a summary of the key findings in section IV.
@@
II Gas heating by individual source
II.1 Leo T profile
The gas in the inner region of Leo T (region I) is predominantly atomic hydrogen, while the gas outside is highly ionized Faerman2013APJ . Given that free electrons in the ionized region cool very efficiently Smith2017MNRAS , we focus on the gas heating in region I. The radius of region I is $r_{\rm I}\simeq 0.35$ kpc. In region I, the average ${\rm H_{I}}$ gas density is $n_{\rm H}=0.07$ ${\rm cm^{-3}}$, electron density is $n_{e}=0.001$ ${\rm cm^{-3}}$, and the DM energy density is $\rho_{\rm DM}=1.75$ GeV ${\rm cm^{-3}}$ Simon2007APJ ; Faerman2013APJ ; Wadekar2021PRD ; Laha2021PLB . The gas column density is $m_{\rm H}n_{\rm H}r_{\rm sys}=1.265\times 10^{-4}$ g ${\rm cm}^{-2}$, where $m_{\rm H}$ is the mass of neutral hydrogen and $r_{\rm sys}=r_{\rm I}$ is the size of the region. The velocity dispersion of the neutral hydrogen (${\rm H_{I}}$) gas is approximately $\sigma_{g}=6.9$ km/s Ryan-Weber2008MNRAS ; Simon2007APJ ; Adams2018AA , leading to an estimated gas temperature of around $T=6000$ K in the region I.
We use the following approximate formula to calculate the gas cooling rate Laha2021PLB ; Wadekar2023PRD ; Smith2017MNRAS
@@ II.1 Leo T profile
where $[{\rm Fe}]/[{\rm H}]=\log_{10}(n_{\rm Fe})/(n_{\rm H})_{\rm gas}$ $\log_{10}(n_{\rm Fe})/(n_{\rm H})_{\rm Sun}$ = $-1.74$ Kirby2013APJ represents the metallicity relative to the Sun, and $\Lambda(T)=2.51\times 10^{-28}T^{0.6}$ is the cooling function valid for $300~{}{\rm K}<T<8000~{}{\rm K}$ Wadekar2021PRD . For this study, we adopt the value
@@
as cited in Lu2021APJL ; Laha2021PLB .
We will use $n_{\rm He}/n_{\rm H}=0.08$ Kirby2013APJ ; Wadekar2021PRD to calculate the interactions of millicharged particles with the interstellar medium gas.
II.2 PBH constraints
@@
where $\rho_{\rm PBH}$ represents the energy density of PBHs. Current constraints on this ratio are approximately (see FIG.10 in Ref.Carr2021RPP and FIG.2 in Ref. Laha2021PLB )
@@
for $M_{\rm PBH}\lesssim 10^{17}$ g, where $M_{\rm PBH}$ denotes the initial mass of PBH. In this paper, the PBHs with $M_{\rm PBH}\lesssim 10^{17}$ g are referred to as the light PBHs. It is known that all DMs could be composed of light PBHs only if $M_{\rm PBH}\simeq 1\times 10^{17}~{}{\rm g}$ Carr2021RPP ; Laha2021PLB . In other mass ranges, light PBHs would constitute only a small fraction of DM Carr2021RPP ; Laha2021PLB . For light PBHs, no accretion disk forms, and only Hawking radiation is available for heating the gas Kim2021MNRAS ; Laha2021PLB ; Melikhov2023PRD .
In contrast, for PBH with $M_{\rm PBH}/M_{\odot}\gtrsim 0.1$, we classify these as heavy PBH, where $M_{\odot}=1.99\times 10^{33}$ g is the solar mass. For such PBHs, Hawking radiation becomes negligible and heat transport from the accretion disk is the primary mechanism for gas heating Lu2021APJL ; Takhistov2022JCAP . The allowed ratio of $f_{\rm PBH}$ for heavy PBHs is approximately (see FIG.10 in Ref.Carr2021RPP and FIG.5 in Ref.Takhistov2022JCAP )
@@
For PBH with intermediate masses, neither Hawking radiation nor heat transport from the accretion disk significantly impacts gas heating.
@@
II.3 Gas heating by light PBHs
@@
For light PBHs, no accretion disk forms, and thus only Hawking radiation contributes to gas heating Kim2021MNRAS ; Laha2021PLB ; Melikhov2023PRD . The heating rate per unit time due to particle $i=(\gamma,e^{\pm})$ emission via Hawking radiation from a single PBH is given by Laha2021PLB
$H_{i{\rm PBH}}=$ $\int_{0}^{\infty}f_{h}(E)(E-m_{i})\frac{d^{2}N_{i}(M_{\rm PBH},a_{\ast})}{dtdE}$
×$\times\left(1-e^{-\tau_{i}}\right)dE,$ (6)
where $m_{i}$ is the mass of particle $i$, $a_{*}=J_{\rm PBH}/(GM_{\rm PBH}^{2})$ is the reduced spin Kerr parameter (with $G$ being the gravitational constant and $J_{\rm PBH}$ the angular momentum of the PBH), and $f_{h}(E)=0.367+0.395(11{\rm eV}/(E-m_{i}))^{0.7}$ represents the fraction of energy loss deposited as heat Kim2021MNRAS ; Laha2021PLB ; Shull1985APJ ; Ricotti2002APJ ; Furlanetto2010MNRAS . The PBH emission spectrum, denotes the number of particles $N_{i}$ emitted per unit energy per unit time, is given by
@@
where $\Gamma$ represents the graybody factor and $E^{\prime}$ denotes the total energy of a particle around rotating PBH. The sign ‘+’ applies to Fermions, while ‘-’ applies to Bosons. The temperature of a charge-neutral rotating PBH is expressed as
@@
In this study, the PBH emission spectra for $\gamma,e^{\pm}$ are generated numerically using the BlackHawk code Arbey2019EPJC . The optical depths of gas for $\gamma$ and $e^{\pm}$ are given by
@@
where $\lambda$ and $S(E)$ represent the absorption length and the stopping power, respectively. We utilize the cumulative photon absorption length from Ref. PDG and the electron stopping power in hydrogen gas from the NIST database Berger2007 .
@@
Light PBHs could constitute all DMs in the very narrow mass range around $M_{\rm PBH}\simeq 1\times 10^{17}~{}{\rm g}$. In other mass ranges, light PBHs can only represent a small fraction of the DM Carr2021RPP ; Laha2021PLB . Therefore, considering the current constraints of $f_{\rm PBH}$, the actual heating ratio in Leo T must be lower than what is shown in FIG. 1. For example, with the constraint $0.05\lesssim f_{\rm PBH}\lesssim 1$, only the mass region $4\times 10^{16}~{}{\rm g}\lesssim M_{\rm PBH}\lesssim 1\times 10^{17}~{}{\rm g}$ in green band in the FIG.1 is relevant. Even using the overestimated curve in FIG. 1, the ratio $Q/C$ falls within small as $1.8\times 10^{-3}\lesssim Q/C\lesssim 8.5\times 10^{-2}$ for $a_{*}=0$ (Schwarzschild BH) and $1.6\times 10^{-2}\lesssim Q/C\lesssim 2.4\times 10^{-1}$ for $a_{*}=0.9$ (rotating BH).
Since $Q_{\rm PBH,HR}<C$ is always satisfied (in other words, $Q_{\rm PBH,HR}$ cannot be comparable to $C$) in this mass range, it is challenging stronger constraints on light PBHs than those previously claimed Ref. Laha2021PLB .
II.4 Gas heating by heavy PBHs
In this subsection, we consider that the heavy PBHs are the sole heating source for the gas in Leo T. For heavy PBHs, Hawking radiation is negligible, and heat transport from the accretion disk becomes the primary mechanism for gas heating Lu2021APJL ; Takhistov2022JCAP . In this case, the three main mechanisms for gas heating are accretion photon emission, dynamical friction, and accretion mass outflow. We assume all heavy PBHs are Schwarzschild black holes and employ the methods and numerical values for model parameters as described in Model 2 of Ref.Takhistov2022JCAP . For the mass range $0.1\leq M_{\rm PBH}/M_{\odot}\leq 30$ (including which is the primary focus of this study), accretion mass outflow is the dominant heating mechanism Takhistov2022JCAP .
Accretion mass outflows, primarily composed protons, become significant in the advection-dominated accretion flow regime within the accretion disk. The heat deposited by these protons is limited by the electronic stopping cross section for protons passing through hydrogen, denoted as $S(E)$, which we obtain from FIG.9 in Ref.Bailey2019PRA . The energy deposited per proton is given by

### v2

@@ 1 Introduction
Dark matter (DM), the enigmatic source of gravity in the universe, remains one of the fundamental challenges in particle physics and cosmology [1]. Among the most compelling candidates for DM is the primordial black hole (PBH) [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13], which formed during the early universe. While PBHs lose mass through the emission of particles via Hawking radiation [14], those with an initial mass exceeding $10^{15}$ g can survive in the current universe without fully evaporating. Although it is unlikely that all DM is composed solely of PBHs across the broad mass spectrum [4], the possibility of PBH formation in numerous early universe models suggest it is reasonable to consider at least a portion of DM consists of PBHs.
@@
If only a portion of DM consists of PBHs, what constitutes the remainder? Since DM has eluded detection through electromagnetic observations, it is typically assumed to lack electromagnetic interactions. Under this assumption, electrically neutral, massive, and stable particles predicted by theories beyond the standard model are considered promising DM candidates [1]. However, some particle DM models propose weak electromagnetic interactions. For example, dark photons may couple very weakly to electrically charged particles via kinetic mixing with ordinary photon [15]. If dark photons is massive, they could serve as DM. Millicharged particles are another potential DM candidate [16]. Additionally, the temperature at cosmic dawn being only half of the expected value [17] can be explained by a Coulomb-like interaction between millicharged particles and neutral hydrogen (${\rm H_{I}}$) [18].
@@
In recent years, the Leo T dwarf galaxy [19], along with other astrophysical gas clouds, has been used to study the heating of interstellar medium gas by PBHs and DM particles. Leo T, located in the Leo constellation, is classified as a transitional object between dwarf spheroidal and dwarf irregular galaxies. Unlike dwarf galaxies within the Milky Way’s virial radius, which often lack gas due to ram pressure stripping [20, 21], Leo T, being beyond this radius, can retain significant amounts of gas. When DMs (such as PBHs, dark photons, or millicharged particles) heat the interstellar medium, the heating rate $Q$ must be equal or less than the astrophysical cooling rate $C$, $Q\leq C$. If $Q$ exceeds $C$, the temperature of the interstellar medium gas would increase, potentially altering the ionization ratio of the gas [22, 25, 23, 28, 26, 27, 24, 15]. In the literature, the heating rate and the cooling rate are depicted as $\dot{Q}$ and $\dot{C}$, respectively; however, we omit ‘dot’ in this paper.
@@
Previously, various sources have been considered for heating interstellar medium gas, including PBHs via Hawking radiation [22, 23, 24], PBHs via accretion disks [25, 26, 27], compact DM [27], dark photons (hidden photons) [15, 31, 29, 28, 30], millicharged particles [32, 28], vector portal DM [31], Higgs portal DM [33], sterile neutrinos [33], dark baryons [33], axions and axion like particles [27, 33], excited DM state [33], magnetically charged DM [27], composite DM [29, 27], annihilating DM [34], electromagnetic dipole moment of DM [35], and DM-baryon interactions [31, 36]. These studies assume that only one type of DM is responsible for heating the interstellar medium gas. For example, in Refs [22, 23, 24, 25, 26], PBHs are considered to be part of the DM and responsible for heating the gas, while other DM are not. Similarly, Ref. [32] assumes that millicharged particles are the sole contributors to the heating of the gas, with other DM having no effect.
@@
In this study, we consider DM to consist of two components: PBHs and additional DM particles. Unlike previous studies [22, 23, 24, 25, 26], which assume that only one type of DM particles heats the gas, we propose that both PBHs and DM particles contribute to the heating of the gas simultaneously in the Leo T dwarf galaxy. We specifically consider dark photons or millichaged particles as the DM particles  [^2] The axions and axion like particles are also good candidate of DM in modern situation [27, 33]. The gas heating via axions and PBHs simultaneously should be investigate. We would like to intend to study this issue in more detail in the future.. The heating mechanism for PBHs is either through Hawking radiation or heat transport from the accretion disk. In contrast, the DM particles heat the gas via mechanism specific to their particle models.
@@
The remainder of this article is organized as follows. Section 2 provides a brief review of the Leo T profile and the current constraints on PBHs. Before exploring the simultaneous heating of the gas in Leo T by both PBHs and DM particles, we first estimate the heating contribution from individual sources, including light PBHs, heavy PBHs, dark photons, and millichaged particles. The main discussion in section 2 are based on Ref. [23] for light PBHs, Refs.[25, 26] for heavy PBHs, and Ref. [28] for dark photons and millicharged particles. In section 3, we assess the combined effect of PBHs and DM particles (dark photons or millicharged particles) on gas heating. We discuss constraints on the PBH fraction, the kinetic mixing parameter of dark photons, and the charge parameter for millicharged particles. The article concludes with a summary of the key findings in section 4.
@@
2 Gas heating by individual source
2.1 Leo T profile
The physical parameters of Leo T are derived from the measured density and velocity dispersion of ${\rm H_{I}}$ gas [38, 37] using an appropriate model. In this study, we adopt the model of Leo T in Ref. [38].
The gas in the inner region of Leo T (region I) is predominantly atomic hydrogen, while the gas outside is highly ionized [38]. Given that free electrons in the ionized region cool very efficiently [39], we focus on the gas heating in region I. The radius of region I is $r_{\rm I}\simeq 0.35$ kpc. In region I, the average ${\rm H_{I}}$ gas density is $n_{\rm H}=0.07$ ${\rm cm^{-3}}$, electron density is $n_{e}=0.001$ ${\rm cm^{-3}}$, and the DM energy density is $\rho_{\rm DM}=1.75$ GeV ${\rm cm^{-3}}$ [40, 38, 28, 23]. The gas column density is $m_{\rm H}n_{\rm H}r_{\rm sys}=1.265\times 10^{-4}$ g ${\rm cm}^{-2}$, where $m_{\rm H}$ is the mass of neutral hydrogen and $r_{\rm sys}=r_{\rm I}$ is the size of the region. The velocity dispersion of the neutral hydrogen (${\rm H_{I}}$) gas is approximately $\sigma_{g}=6.9$ km/s [37, 40, 41], leading to an estimated gas temperature of around $T=6000$ K in the region I.
We use the following approximate formula to calculate the gas cooling rate [23, 27, 39]
@@ 2.1 Leo T profile
where $[{\rm Fe}]/[{\rm H}]=\log_{10}(n_{\rm Fe})/(n_{\rm H})_{\rm gas}$ $\log_{10}(n_{\rm Fe})/(n_{\rm H})_{\rm Sun}$ = $-1.74$ represents the metallicity relative to the Sun[42], and $\Lambda(T)=2.51\times 10^{-28}T^{0.6}$ is the cooling function valid for $300~{}{\rm K}<T<8000~{}{\rm K}$ [28]. For this study, we adopt the value
@@
as cited in Refs. [25, 23].  [^3] In the region I, the ${\rm H_{I}}$ density is approximately constant [28]. For this reason, the gas temperature and the cooling rate are assumed to be constant in this study.
We will use $n_{\rm He}/n_{\rm H}=0.08$ to calculate the interactions of millicharged particles with the interstellar medium gas [42, 28].
2.2 PBH constraints
@@
where $\rho_{\rm PBH}$ represents the energy density of PBHs. Current constraints on this ratio are approximately (see FIG.10 in Ref.[4] and FIG.2 in Ref. [23])
@@
for $M_{\rm PBH}\lesssim 10^{17}$ g, where $M_{\rm PBH}$ denotes the initial mass of PBH. The first line in Eq. (4) seems to have to be dependent from mass of PBH; however, it is just an approximation. In this paper, the PBHs with $M_{\rm PBH}\lesssim 10^{17}$ g are referred to as the light PBHs. It is known that all DMs could be composed of light PBHs only if $M_{\rm PBH}\simeq 1\times 10^{17}~{}{\rm g}$ [4, 23]. In other mass ranges, light PBHs would constitute only a small fraction of DM [4, 23]. For light PBHs, no accretion disk forms, and only Hawking radiation is available for heating the gas [22, 23, 24].
In contrast, for PBH with $M_{\rm PBH}/M_{\odot}\gtrsim 0.1$, we classify these as heavy PBH, where $M_{\odot}=1.99\times 10^{33}$ g is the solar mass. For such PBHs, Hawking radiation becomes negligible and heat transport from the accretion disk is the primary mechanism for gas heating [25, 26]. The allowed ratio of $f_{\rm PBH}$ for heavy PBHs is approximately (see FIG.10 in Ref.[4] and FIG.5 in Ref.[26])
@@
For PBH with intermediate masses, $10^{17}{\rm g}\lesssim M_{\rm PBH}\lesssim 0.1M_{\odot}$, neither Hawking radiation nor heat transport from the accretion disk significantly impacts gas heating.
@@
2.3 Gas heating by light PBHs
@@
For light PBHs, no accretion disk forms, and thus only Hawking radiation contributes to gas heating [22, 23, 24]. The heating rate per unit time due to particle $i=(\gamma,e^{\pm})$ emission via Hawking radiation from a single PBH is given by [23]
[…]
where $m_{i}$ is the mass of particle $i$, $a_{*}=J_{\rm PBH}/(GM_{\rm PBH}^{2})$ is the reduced spin Kerr parameter (with $G$ being the gravitational constant and $J_{\rm PBH}$ the angular momentum of the PBH), and $f_{h}(E)=0.367+0.395(11{\rm eV}/(E-m_{i}))^{0.7}$ represents the fraction of energy loss deposited as heat [22, 23, 43, 44, 45]. (See also Refs.[46, 47, 48, 49, 50, 51, 52]). The PBH emission spectrum, denotes the number of particles $N_{i}$ emitted per unit energy per unit time, is given by
@@
where $\Gamma$ represents the graybody factor[53] and $E^{\prime}$ denotes the total energy of a particle around rotating PBH. The sign ‘+’ applies to Fermions, while ‘-’ applies to Bosons. The temperature of a charge-neutral rotating PBH is expressed as
@@
In this study, the PBH emission spectra for $\gamma,e^{\pm}$ are generated numerically using the BlackHawk code [54]. The optical depths of gas for $\gamma$ and $e^{\pm}$ are given by
@@
where $\lambda$ and $S(E)$ represent the absorption length in unit of $\rm{g~{}cm^{-2}}$ and the stopping power in unit of $\rm{MeV~{}(g~{}cm^{-2})^{-1}}$, respectively. We utilize the cumulative photon absorption length from Ref. [55] and the electron stopping power in hydrogen gas from the NIST database [56].
@@
Light PBHs could constitute all DMs in the very narrow mass range around $M_{\rm PBH}\simeq 1\times 10^{17}~{}{\rm g}$. In other mass ranges, light PBHs can only represent a small fraction of the DM [4, 23]. Therefore, considering the current constraints of $f_{\rm PBH}$, the actual heating ratio in Leo T must be lower than what is shown in FIG. 1. For example, with the constraint $0.05\lesssim f_{\rm PBH}\lesssim 1$, only the mass region $4\times 10^{16}~{}{\rm g}\lesssim M_{\rm PBH}\lesssim 1\times 10^{17}~{}{\rm g}$ in green band in the FIG. 1 is relevant. Even using the overestimated curve in FIG. 1, the ratio $Q/C$ falls within small as $1.8\times 10^{-3}\lesssim Q/C\lesssim 8.5\times 10^{-2}$ for $a_{*}=0$ (Schwarzschild BH) and $1.6\times 10^{-2}\lesssim Q/C\lesssim 2.4\times 10^{-1}$ for $a_{*}=0.9$ (rotating BH).
Since $Q_{\rm PBH,HR}<C$ is always satisfied (in other words, $Q_{\rm PBH,HR}$ cannot be comparable to $C$) in this mass range, it is challenging stronger constraints on light PBHs than those previously claimed Ref. [23]. We would like to note that the possible suppression of Hawking radiation related to the information paradox has recently been discovered [57, 58, 59, 60]. The inclusion of such effects, e.g. memory burden effect, may change the results of this paper. We would like to keep this issue in our future study.
2.4 Gas heating by heavy PBHs
In this subsection, we consider that the heavy PBHs are the sole heating source for the gas in Leo T. For heavy PBHs, Hawking radiation is negligible, and heat transport from the accretion disk becomes the primary mechanism for gas heating [25, 26]. In this case, the three main mechanisms for gas heating are accretion photon emission, dynamical friction, and accretion mass outflow. We assume all heavy PBHs are Schwarzschild black holes and employ the methods and numerical values for model parameters as described in Model 2 of Ref.[26]. For the mass range $0.1\leq M_{\rm PBH}/M_{\odot}\leq 30$ (including which is the primary focus of this study), accretion mass outflow is the dominant heating mechanism [26].
Accretion mass outflows, primarily composed protons, become significant in the advection-dominated accretion flow regime within the accretion disk. The heat deposited by these protons is limited by the electronic stopping cross section for protons passing through hydrogen, denoted as $S(E)$, which we obtain from FIG.9 in Ref.[61]. The energy deposited per proton is given by